OpenStudy (anonymous):

A corner store sells two kinds of baked goods: cakes and pies. A cake costs $5 and a pie costs $7. In one day, the store sold 15 baked goods for a total of $91. How many cakes did they sell? A. 7 cakes B. 4 cakes C. 8 cakes D. 5 cakes

OpenStudy (anonymous):

can you give me the answer ? it wouldnt show

OpenStudy (anonymous):

5x+7y=91 x+y=15 solve for y in x+y=15 and plug that into y in 5x+7y=91 and solve for x

OpenStudy (anonymous):

how do you solve for x

OpenStudy (anonymous):

have you done the first two steps? (solve for y and plug it into the other equation?)